OpenStudy (anonymous):

Events A and B are overlapping. P(A) = 6/11 P(B) = 4/11 P(A and B) =2/11 Find P(A or B).

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

P(A or B) = 8/11 P(Only A) = 4/11 P(Only B) = 2/11 Are you familiar with Venn Diagrams?

OpenStudy (anonymous):

If your teacher hasn't taught the probability rules, or if the rules are too confusing, Venn Diagrams are lifesavers.

OpenStudy (anonymous):

8/11?